# SemHARQ: Semantic-Aware HARQ for Multi-task Semantic Communications

### v1

@@ Abstract
Intelligent task-oriented semantic communications (SemComs) have witnessed great progress with the development of deep learning (DL). In this paper, we propose a semantic-aware hybrid automatic repeat request (SemHARQ) framework for the robust and efficient transmissions of semantic features. First, to improve the robustness and effectiveness of semantic coding, a multi-task semantic encoder is proposed. Meanwhile, a feature importance ranking (FIR) method is investigated to ensure the important features delivery under limited channel resources. Then, to accurately detect the possible transmission errors, a novel feature distortion evaluation (FDE) network is designed to identify the distortion level of each feature, based on which an efficient HARQ method is proposed. Specifically, the corrupted features are retransmitted, where the remaining channel resources are used for incremental transmissions. The system performance is evaluated under different channel conditions in multi-task scenarios in Internet of Vehicles. Extensive experiments show that the proposed framework outperforms state-of-the-art works by more than $20\%$ in rank-1 accuracy for vehicle re-identification, and $10\%$ in vehicle color classification accuracy in the low signal-to-noise ratio regime.
@@ I Introduction
In the era of Internet of Everything (IoE), communications among humans, machines, objects, and intelligent agents enable intelligent services to individuals and industries. As a result, the requirements for communication systems to effectively convey the intended information are increased. Conventional communication systems focus on the successful transmissions of bits but ignore the meanings of messages, facing critical challenges to support the aforementioned applications over limited communication resources [1, 2, 3]. To solve this issue, semantic communications (SemComs), which transmit the intention-related semantics, can significantly reduce the amount of transmission data while ensuring the system performance. Consequently, the efficiency of communication is improved effectively. Meanwhile, with the development of artificial intelligence (AI), deep learning (DL) has shown its superiority in semantic representation, compression and recovery, which has been widely used in task-oriented SemComs [4, 5].
@@
To guarantee the reliability of semantic transmissions, most of the current DL-based SemCom systems transmit task-related semantics and identify semantic errors by joint source-channel coding (JSCC) [6, 7]. Particularly, Bourtsoulatze et al. propose a JSCC technique for image transmission, which directly maps the image pixel values to the complex-valued channel input symbols [6]. Furthermore, besides transmitting semantics in an end-to-end (E2E) manner, a multi-user cooperative semantic framework based on JSCC is proposed to exploit the correlations among users in [7]. However, the above SemCom systems are designed with a fixed coding rate, incapable of transmitting semantics adaptively according to the dynamic channel conditions. To overcome this problem, a variable length coding method is proposed based on dynamic neural networks in [8]. Meanwhile, a policy network is trained additionally in [9] to select a portion of the encoded feature vector based on the signal-to-noise ratio (SNR). Nevertheless, the robustness of semantic recovery is limited in the above systems due to the lack of precise process of semantic errors.
@@
To improve the error correction ability in communication systems, hybrid automatic repeat request (HARQ) is commonly used in conventional communications [10, 11]. Specifically, a cyclic redundancy check (CRC) error detection technique is combined with retransmissions to improve transmission reliability at the data link layer. Nevertheless, the meanings of information are ignored in the above method, making it unable to handle semantic errors. Therefore, to further enhance the robustness and efficiency of SemComs, semantic retransmission schemes have been studied [12, 13, 14, 15, 16]. In [12], the data sample is thoroughly retransmitted in a DL-based communication network when the received semantic information is not sufficient to complete the task. A channel output feedback is further incorporated into the retransmission mechanism in [13] to improve the reconstruction quality. Although constantly retransmitting messages can improve the performance of specific intelligent task, the efficiency in resource utilization is too low for practical usage. Therefore, incremental transmission methods are further incorporated to improve the transmission efficiency. Specifically, multiple incremental transmissions are supported by different autoencoders in [17] to reduce semantic transmission error. To further simplify the network architecture, a unified single decoder is proposed in [9] to exploit the incremental knowledge (IK) obtained from retransmission. Besides the text and image transmission, an incremental redundancy HARQ framework is investigated for video conferencing with a semantic error detector [18]. However, the above works retransmit the same messages simply or directly transmit incremental redundancy triggered by task performance, having no insight into the specific distortion from the channel fading and noise. Consequently, the error detection on the semantic feature dimension is always ignored, where the features that are successfully received and decoded are probable to be transmitted again, decreasing the efficiency of resource utilization. At the same time, the receiver may fail to recover critical semantics that are previously transmitted when only incremental transmissions are involved without considering retransmissions, degrading the system performance.
@@
    Then, to further improve the transmission efficiency, a feature importance ranking (FIR) method is designed for multi-task SemComs. Specifically, a multi-task semantic encoder is proposed [23, 24, 25, 26], where the correlated semantics among different tasks are extracted jointly, making the semantic coding more robust and efficient than existing single-task coding [27, 28, 29, 30]. Moreover, the semantic importance levels are further evaluated by measuring the sensitivity of task performance to different features, where the gradients of the task performance w.r.t. the features are considered [31].
@@ II System Model
In this section, the system model of SemHARQ is introduced. Specifically, $K$ intelligent tasks are executed simultaneously in intelligent transportation scenarios, and the inputs of the system are assumed to be vehicle images, as shown in Fig. 1. Note that the system can be readily extended to other scenarios and modalities. First, the semantic features are extracted and then transmitted to the receiver after scalable selection according to the importance levels. Then, the SemHARQ is triggered after receiving the semantic features from the channel. After all the transmissions, the concatenated feature vector is decoded by a joint source-channel decoder (JSC-Decoder) and multiple semantic decoders. Finally, the multiple tasks are performed by different task performers.
II-A Transmitter
As shown in Fig. 2, the input images denoted as $\bm{x}\in\mathbb{R}^{\widetilde{B}\times W\times H}$ are encoded by a multi-task semantic encoder, where $\widetilde{B}$ is the batch size, and $W$, $H$ are the width and height of an image, respectively. The semantic features are extracted as $\bm{s}\in\mathbb{R}^{N}$, where $N$ is the number of features in $\bm{s}$. Then, a JSC-Encoder encodes $\bm{s}$ as $\bm{f}\in\mathbb{C}^{L}$, where $L$ is the length of the feature vector. Afterwards, the feature vector selected by a scalable feature selector for HARQ is represented as ${\bm{\widetilde{z}}}_{j}\in\mathbb{C}^{B}$, where $B$ is the number of features to be transmitted according to channel state information (CSI), $j$ is the number of retransmissions. The value of $B$ is determined by the transmission bandwidth in a single time slot so that the length of the transmitted feature vector can be scalably adjusted according to the channel conditions. Finally, the average power control is utilized to normalize the input signals of the channel at each transmission, which is shown as
${\bm{z}}_{j}=\sqrt{PB}\frac{{\bm{\widetilde{z}}}_{j}}{||{\bm{\widetilde{z}}}_{j}||_{2}},$ (1)
@@ II-A Transmitter
Furthermore, ${\bm{z}}_{j}$ is sent and passes through a stochastic physical channel. The channel output can be represented as
${\bm{\hat{z}}}_{j}=\bm{h}{\bm{z}}_{j}+\bm{n},$ (2)
where $\bm{n}\in\mathbb{C}^{B}$ denotes the white Gaussian noise, the value of which follows $\bm{n}_{i}\sim\mathcal{C}\mathcal{N}(0,\sigma^{2})$. $\bm{h}\in\mathbb{C}^{B}$ represents the fading coefficient of the channel. In this paper, the Rician channel is considered. The fading coefficient follows $\mathcal{C}\mathcal{N}\Bigl{(}\sqrt{r/(r+1)},1/(r+1)\Bigl{)}$, where $r$ is set to $2$.
II-B Receiver
At the receiver, ${\bm{\hat{z}}}_{j}$ is measured by a retransmission identification module to detect if the retransmission is required. If ${\bm{\hat{z}}}_{j}$ passes the detection, it will be directly concatenated with the previously received features and the concatenated vector is denoted as ${\bm{\overline{z}}}_{j}=[{\bm{\hat{z}}}_{1},...,{\bm{\hat{z}}}_{j}]$. Otherwise, the received ${\bm{\hat{z}}}_{j}$ will be evaluated by the FDE network to generate a binary feedback vector $\bm{p}_{j+1}$, indicating which features in ${\bm{\hat{z}}}_{j}$ are needed to be retransmitted. Such retransmissions will be repeated until the success of multi-task or the maximum times of retransmission is reached. After all the transmissions, the concatenated vector is decoded as $\bm{\hat{s}}\in\mathbb{R}^{N}$ by the JSC-Decoder. $\bm{\hat{s}}$ is then recovered to $[\bm{\hat{y}}_{1},\bm{\hat{y}}_{2},...,\bm{\hat{y}}_{K}]$ by multiple semantic decoders, where $K$ represents the number of tasks. Finally, the task prediction vectors $[\bm{c}_{1},\bm{c}_{2},...,\bm{c}_{K}]$ are generated from the task performers.
@@ III Feature distortion evaluation and importance ranking
In this section, the FDE network and FIR method are introduced. Firstly, the FDE network is investigated to locate the corrupted features that need to be retransmitted in the received feature vector. Then, the FIR method is designed to identify important features in the scalable feature selector for HARQ.
III-A Feature distortion evaluation (FDE) network
Figure 3: The structure of the feature distortion evaluation (FDE) network.
@@ III-A Feature distortion evaluation (FDE) network
As shown in Fig. 3, to obtain the distortion extents of the received features, a distortion measure module $\mathcal{F}$ is proposed. Specifically, a neural network (NN) consisting of four dense layers is utilized to evaluate the distortion distribution of the received vector ${\bm{\hat{z}}}_{j}$ after $j$ retransmissions. On this basis, a distortion measure vector is generated by $\mathcal{F}$ during the E2E training, which has the same dimension as ${\bm{\hat{z}}}_{j}$ and indicates the distortion levels of all the features in ${\bm{\hat{z}}}_{j}$.
@@
To enable $\mathcal{F}$ to distinguish the correlations between input features and the undistorted feature vector $\bm{s}$, thus inferring the distortion extents of these features, $\mathcal{F}$ is trained independently before the entire system training. Since mutual information can be used to measure the semantic similarity between the transmitted and received features, the mutual information of $\bm{s}$ and the JSC-decoded $\bm{\hat{s}}$ is estimated and serves as the training objective of $\mathcal{F}$. Consequently, with the parameters optimized by the mutual information, the contributions of different input features to the information recovery can be quantified by $\mathcal{F}$. The optimization problem is defined as
$\max_{\xi_{\mathcal{F}}}\;\;I(\bm{s};\bm{\hat{s}}),$ (3)
where $\xi_{\mathcal{F}}$ is the parameter set of $\mathcal{F}$. Specifically, the mutual information of $\bm{s}$ and $\bm{\hat{s}}$ can be calculated by
$I(\bm{s};\bm{\hat{s}})={\mathbb{E}}_{p(s,\hat{s})}\bigg{[}\log\frac{p(s,\hat{s})}{p(s)p(\hat{s})}\bigg{]}=D_{\rm KL}\Bigl{(}p(s,\hat{s})||p(s)p(\hat{s})\Bigl{)},$ (4)
@@
where $p(s)$ and $p(\hat{s})$ are the marginal probability of $\bm{s}$ and $\bm{\hat{s}}$, respectively, and $p(s,\hat{s})$ is the joint probability of $\bm{s}$ and $\bm{\hat{s}}$. Meanwhile, the mutual information is equivalent to the Kullback-Leibler (KL) divergence between the product of the marginal probabilities and the joint probability. Inspired by the general-purpose parametric neural estimation method of mutual information [32], the KL-divergence in (4) obeys the following dual representation,
$D_{\rm KL}\Bigl{(}p(s,\hat{s})||p(s)p(\hat{s})\Bigl{)}=$ (5)
$\sup_{T:\Omega\rightarrow\mathbb{R}}\mathbb{E}_{p(s,\hat{s})}[T]-\log\left(\mathbb{E}_{p(s)p(\hat{s})}\left[e^{T}\right]\right),$
where the supremum is taken over all functions $T$. Consequently, the lower bound of $I(\bm{s};\bm{\hat{s}})$ can be derived as follows,
$I(\bm{s};\bm{\hat{s}})$ $=D_{\rm KL}\Bigl{(}p(s,\hat{s})||p(s)p(\hat{s})\Bigl{)}$ (6)
$\geqslant{\mathbb{E}}_{p\left({s,\hat{s}}\right)}\left[T\right]-\log\left({{{\mathbb{E}}_{p\left(s\right)p\left(\hat{s}\right)}}[{e^{T}}]}\right).$
@@
Due to $T$ can be any class of functions satisfying integrability, the NN in $\mathcal{F}$ is employed as $T$ to obtain a precise bound according to (6). Finally, the estimated mutual information can be denoted as
@@
where $\mathcal{F}$ is trained in an unsupervised manner and the optimization goal is updated to maximize $I(\bm{s};\bm{\hat{s}}|\mathcal{F})$. Additionally, note that during the training phase of $\mathcal{F}$, the inputs are generated by sampling from $(\bm{s},\bm{\hat{s}})$, $\bm{s}$, and $\bm{\hat{s}}$. Moreover, an output vector ${\overline{\bm{{d}}}}_{j}$ is obtained by feeding ${\bm{\hat{z}}}_{j}$ into $\mathcal{F}$ when the entire SemHARQ system is trained, and the distortion measure vector ${\bm{{d}}}_{j}$ can be given by ${\bm{{d}}}_{j}=1-{\overline{\bm{{d}}}}_{j}$.
@@
Furthermore, to instruct the transmitter to retransmit the corrupted features, ${\bm{{d}}}_{j}$ is quantized into a binary feedback vector ${\bm{{p}}}_{j+1}$ to trigger the $(j+1)$-th retransmission. Note that the process is not trivial because the quantization is discrete, which makes the network non-differentiable and hard to optimize with back-propagation. Thus, a straight-through Gumbel-softmax estimator is investigated [33], where ${\bm{{d}}}_{j}$ is discretized in the forward pass but a gradient estimator based on Gumbel-softmax distributions is employed to resolve the non-differentiability in the backward pass. Specifically, for the forward propagation, ${\bm{{p}}}_{j+1}$ is drawn from ${\bm{{d}}}_{j}$ using a distortion threshold $t$, which can be represented as
${p}_{(j+1)i}=\left\{\begin{array}[]{ll}1&\text{ if }{d}_{ji}\geq t,\\ 0&\text{ otherwise, }\end{array}\right.$ (8)
@@
where ${p}_{(j+1)i}$ and ${d}_{ji}$ denote the $i$-th value of ${\bm{{p}}_{j+1}}$ and ${\bm{{d}}_{j}}$, respectively. On this basis, ${\bm{{p}}_{j+1}}$ can serve as a feedback vector to guide the next retransmission. Particularly, the number of features requested to be retransmitted is equal to the number of $1$ values in ${\bm{{p}}_{j+1}}$, which is given by
$R=\sum_{i=1}^{B}{p}_{(j+1)i},$ (9)
where $B$ is the length of ${\bm{{p}}_{j+1}}$.

### v2

# SemHARQ: Semantic-Aware Hybrid Automatic Repeat Request for Multi-Task Semantic Communications

@@ Abstract
Intelligent task-oriented semantic communications (SemComs) have witnessed great progress with the development of deep learning (DL), where multi-task SemComs that perform multiple tasks simultaneously attach great importance due to its high efficiency. However, the study of robust multi-task-oriented semantics transmission is still in early stages. In this paper, we propose a semantic-aware hybrid automatic repeat request (SemHARQ) framework for the robust and efficient transmissions of multi-task semantic features. First, to improve the robustness and effectiveness of semantic coding, a multi-task semantic encoder is proposed. Meanwhile, a feature importance ranking (FIR) method is investigated to ensure the important features delivery under limited channel resources. Then, to accurately detect the possible transmission errors, a novel feature distortion evaluation (FDE) network is designed to identify the distortion level of each feature, based on which an efficient HARQ method is proposed. Specifically, the corrupted features are retransmitted, where the remaining channel resources are used for incremental transmissions. The system performance is evaluated under different channel conditions in multi-task scenarios in the Internet of Vehicles. Extensive experiments show that the proposed framework outperforms state-of-the-art works by more than $20\%$ in rank-1 accuracy for vehicle re-identification, and $10\%$ in vehicle color classification accuracy in the low signal-to-noise ratio regime.
@@ I Introduction
In the era of the sixth-generation (6G) wireless communications, communications among humans, machines, objects, and intelligent agents enable intelligent services to individuals and industries [1, 2, 3]. As a result, the requirements for communication systems to effectively convey the intended meaning have increased. Conventional communication systems focus on the successful transmissions of bits but ignore the meanings of messages, facing critical challenges to support the aforementioned applications over limited communication resources [4, 5]. To solve this issue, semantic communications (SemComs), which transmit the intention-related semantics, can significantly reduce the amount of transmission data while ensuring the system performance. Consequently, the efficiency of communication is improved effectively. Meanwhile, with the development of artificial intelligence (AI), deep learning (DL) has shown its superiority in semantic representation [6], compression [7] and recovery [8], which has been widely used in task-oriented SemComs to enable the extraction, compression, and reconstruction of semantic features [9, 10].
@@
To guarantee the reliability of semantic transmissions, most of the current DL-based SemCom systems transmit task-related semantics and identify semantic errors by joint source-channel coding (JSCC) [11, 12]. Particularly, Bourtsoulatze et al. propose a JSCC technique for image transmission, which directly maps the image pixel values to the complex-valued channel input symbols [11]. Furthermore, besides transmitting semantics in an end-to-end (E2E) manner, a multi-user cooperative semantic framework based on JSCC is proposed to exploit the correlations among users in [12]. However, the above SemCom systems are designed with a fixed coding rate, incapable of transmitting semantics adaptively according to the dynamic channel conditions. To overcome this problem, a variable length coding method is proposed based on dynamic neural networks in [13]. Meanwhile, a policy network is trained additionally in [14] to select a part of the encoded feature vector based on the signal-to-noise ratio (SNR). Nevertheless, the robustness of semantic recovery is limited in the mentioned systems due to the lack of an accurate processing on semantic errors.
To improve the robustness of communication systems, hybrid automatic repeat request (HARQ) is commonly used in conventional communications [15]. Specifically, the received information packet is detected by a cyclic redundancy check (CRC) detector at the receiver. Then, a NACK signal is sent back to the transmitter to trigger a retransmission if the detection fails. The process will repeat until the packet is received correctly, or the maximum number of transmissions is reached. Particularly, based on the type of retransmitted data, the conventional HARQ can be classified into three main types. In type-I HARQ (HARQ-I), the same data packet is retransmitted, and the erroneous packet is replaced with the new one [16]. Since the erroneous packets are simply discarded in HARQ-I, the useful information is not fully utilized, reducing the efficiency of retransmissions. To overcome this problem, in type-II HARQ (HARQ-II), the original bit vector is punctured into different lengths, and the incremental redundancy is transmitted when retransmission is needed. Then, at the receiver, the received packet is combined with the previously transmitted information for the forward error correction (FEC) decoding [17]. However, due to the lack of information bits, the received incremental redundancy cannot be decoded independently, degrading the efficiency of decoding. Therefore, type-III HARQ (HARQ-III) is proposed to include the information bits and parity bits simultaneously at every retransmission [18]. Nevertheless, the meanings of transmitted information are ignored in the above methods, making it incapable of handling semantic errors. Moreover, since the CRC detector can only detect whether the entire received packet is erroneous, the fine-grained error detection is not considered, degrading the efficiency of error detection.
@@
Therefore, to further enhance the robustness and efficiency of SemComs, semantic retransmission schemes have been studied [19, 20, 21, 22, 23, 24]. In [19], the data sample is thoroughly retransmitted in a DL-based communication network when the received semantic information is not sufficient to complete the task. A channel output feedback is further incorporated into the retransmission mechanism in [20] to improve the reconstruction quality. Although constantly retransmitting messages can improve the performance of specific intelligent task, the efficiency in resource utilization is too low for practical usage. Therefore, incremental transmission methods are further incorporated to improve the transmission efficiency. Specifically, multiple incremental transmissions are supported by different autoencoders in [25] to reduce semantic transmission error. To further simplify the network architecture, a unified single decoder is proposed in [14] to exploit the incremental knowledge (IK) obtained from retransmission. Besides the text and image transmission, an incremental redundancy HARQ framework is investigated for video conferencing with a semantic error detector [26]. However, the above works retransmit the same messages simply or directly transmit incremental redundancy triggered by task performance, having no insight into the specific distortion from the channel fading and noise. Consequently, the error detection on the semantic feature dimension is always ignored, where the features that are successfully received and decoded are probable to be transmitted again, decreasing the efficiency of resource utilization. At the same time, the receiver may fail to recover critical semantics that are previously transmitted when only incremental transmissions are involved without considering retransmissions, degrading the system performance.
@@
    Then, to further improve the transmission efficiency, a feature importance ranking (FIR) method is designed for multi-task SemComs. Specifically, a multi-task semantic encoder is proposed [31, 32, 33], where the correlated semantics among different tasks are extracted jointly, making the semantic coding more robust and efficient than existing single-task coding [34, 35]. Moreover, the semantic importance levels are further evaluated by measuring the sensitivity of task performance to different features, where the gradients of the task performance with respect to the features are considered [36].
@@
Notation: $\mathbb{C}^{n}$ and $\mathbb{R}^{n}$ represent the set of $n$-dimensional complex and real vectors, respectively. Bold-font variables denote vectors, where $x_{i}$ represents the $i$-th value of vector $\bm{x}$. Additionally, note that in this paper, the superscript of a vector indicates the number of retransmissions. For example, $\bm{x}^{j}$ indicates the vector corresponding to the $j^{th}$ retransmission. Finally, $x\sim\mathcal{CN}(\mu,\sigma^{2})$ means variable $x$ follows a circularly-symmetric complex Gaussian distribution with mean $\mu$ and covariance $\sigma^{2}$.
@@ II System Model
In this section, the system model of SemHARQ is introduced. Specifically, $K$ intelligent tasks are executed simultaneously in intelligent transportation scenarios, and the inputs of the system are assumed to be vehicle images, as shown in Fig. 1. Note that the system can be readily extended to other scenarios and modalities. First, the semantic features are extracted and encoded as symbols, then directly transmitted to the receiver after mapping into complex symbols and scalably selected according to the importance levels. Then, the SemHARQ is triggered after receiving the semantic features from the channel. After all the transmissions, the concatenated feature vector is decoded by a joint source-channel decoder (JSC-Decoder) and multiple semantic decoders. Finally, the multiple tasks are performed by different task performers.
As shown in Fig. 2, the input images denoted as $\bm{x}\in\mathbb{R}^{\widetilde{B}\times W\times H}$ are encoded by a multi-task semantic encoder, where $\widetilde{B}$ is the batch size, and $W$, $H$ are the width and height of an image, respectively. The semantic features are extracted as $\bm{s}\in\mathbb{R}^{N}$, where $N$ is the number of features in $\bm{s}$. Then, a JSC-Encoder encodes $\bm{s}$ as $\bm{f}\in\mathbb{C}^{L}$, where $L$ is the length of the feature vector. Afterwards, the feature vector selected by a scalable feature selector for HARQ is represented as ${\bm{\widetilde{z}}}^{j}\in\mathbb{C}^{B}$, where $B$ is the number of features to be transmitted according to channel state information (CSI), $j$ is the number of retransmissions. The value of $B$ is determined by the transmission bandwidth in a single time slot so that the length of the transmitted feature vector can be scalably adjusted according to the channel conditions. Finally, the average power control is utilized to normalize the input signals of the channel at each transmission, which is shown as
[…]
@@
Furthermore, ${\bm{z}}^{j}$ is sent and passes through the stochastic physical channel. The channel output can be represented as
${\bm{\hat{z}}}^{j}=\bm{h}{\bm{z}}^{j}+\bm{n},$ (2)
where $\bm{n}\in\mathbb{C}^{B}$ denotes the white Gaussian noise, the value of which follows ${n}_{i}\sim\mathcal{C}\mathcal{N}(0,\sigma^{2})$. $\bm{h}\in\mathbb{C}^{B}$ represents the fading coefficient of the channel. In this paper, the performance is evaluated in the intelligent transportation scenario as shown in Fig. 1, where both the line-of-sight (LOS) and the non-line-of-sight (NLOS) paths should be considered [37, 38]. As a result, the Rician channel is considered. The fading coefficient follows $\mathcal{C}\mathcal{N}\Bigl{(}\sqrt{r/(r+1)},1/(r+1)\Bigl{)}$, where $r$ is set to $2$.
At the receiver, ${\bm{\hat{z}}}^{j}$ is measured by a retransmission identification module to detect if the retransmission is required. If ${\bm{\hat{z}}}^{j}$ passes the detection, it will be directly concatenated with the previously received features and the concatenated vector is denoted as ${\bm{\overline{z}}}^{j}=[{\bm{\hat{z}}}^{1},...,{\bm{\hat{z}}}^{j}]$. Otherwise, the received ${\bm{\hat{z}}}^{j}$ will be evaluated by the feature distortion evaluation (FDE) network to generate a binary feedback vector ${\bm{p}}^{(j+1)}$, indicating which features in ${\bm{\hat{z}}}^{j}$ are needed to be retransmitted. Such retransmissions will be repeated until the success of multi-task or the maximum times of retransmission is reached. After all the transmissions, the concatenated vector is decoded as $\bm{\hat{s}}\in\mathbb{R}^{N}$ by the JSC-Decoder. $\bm{\hat{s}}$ is then semantically decoded to $\bm{\hat{y}}_{1},\bm{\hat{y}}_{2},...,\bm{\hat{y}}_{K}]$ by multiple semantic decoders, where $K$ represents the number of tasks. Finally, the task prediction vectors $\bm{c}_{1},\bm{c}_{2},...,\bm{c}_{K}]$ are generated from the task performers.
@@ III Feature Distortion Evaluation and Importance Ranking
In this section, the FDE network and feature importance ranking (FIR) method are introduced. Firstly, the FDE network is investigated to locate the corrupted features that need to be retransmitted in the received feature vector. Then, the FIR method is designed to identify important features in the scalable feature selector for HARQ.
III-A FDE network
@@
As shown in Fig. 3, to obtain the distortion extents of the received features, a distortion measure module $\mathcal{F}$ is proposed. Specifically, a neural network (NN) consisting of four dense layers is utilized to evaluate the distortion distribution of the received vector ${\bm{\hat{z}}}^{j}$ after $j^{th}$ retransmission. Since ${\bm{\hat{z}}}^{j}$ corresponds to a subset of the original features after scalable feature selection, to align the features in ${\bm{\hat{z}}}^{j}$ with the original feature vector $\bm{s}$, a vector ${\bm{\hat{z}}}^{j}_{\mathrm{p}}$ of the same dimension as the original vector is generated from ${\bm{\hat{z}}}^{j}$ by padding the other positions with zeros. Then, a distortion measure vector can be generated by feeding ${\bm{\hat{z}}}^{j}_{\mathrm{p}}$ into $\mathcal{F}$, which has the same dimension as ${\bm{\hat{z}}}^{j}_{\mathrm{p}}$, inferring the distortion levels of all the features in ${\bm{\hat{z}}}^{j}$.
@@
To enable $\mathcal{F}$ to learn the discrepancy of the distributions between the received features corrupted by the channel impairments and the original undistorted feature vector $\bm{s}$, $\mathcal{F}$ is trained independently before the entire system training. Since mutual information can be used to measure the semantic similarity between the transmitted and received features, the mutual information of the original feature vector $\bm{s}$ and the JSC-decoded feature vector $\bm{\hat{s}}$ is estimated and serves as the training objective of $\mathcal{F}$.
Specifically, the optimization problem of FDE network training is defined as
[…]
The mutual information of $\bm{s}$ and $\bm{\hat{s}}$ can be calculated as [39],
$),I(\bm{s};\bm{\hat{s}})={\mathbb{E}}_{p(s,\hat{s})}\bigg{[}\log\frac{p(s,\hat{s})}{p(s)p(\hat{s})}\bigg{]}=D_{\rm KL}\Bigl{(}p(s,\hat{s})||p(s)p(\hat{s}$$\Bigl{)}$, (4)
@@
where $p(s)$ and p(s^)p($\hat${s}) are the marginal probability of $\bm{s}$ and $\bm{\hat{s}}$, respectively, and p(s,s^)p($\hat{s}$) is the joint probability of $\bm{s}$ and $\bm{\hat{s}}$. Note that the mutual information can be represented by the Kullback-Leibler (KL) divergence between the product of the marginal probabilities and the joint probability. Inspired by the general-purpose parametric mutual information neural estimation (MINE) method of  [40], the KL-divergence in Eq. (4) obeys the following dual representation [40, Eq. (5)],
$)=\displaystyle D_{\rm KL}\Bigl{(}p(s,\hat{s})||p(s)p(\hat{s}$$\Bigl{)}$= (5)
$\sup_{\mathcal{T}:\Omega\rightarrow\mathbb{R}}\mathbb{E}_{p(s,\hat{s})}[\mathcal{T}]-\log\left(\mathbb{E}_{p(s)p(\hat{s})}\left[e^{\mathcal{T}}\right]\right),$
[…]
$\geqslant{\mathbb{E}}_{p\left({s,\hat{s}}\right)}\left[\mathcal{T}\right]-\log\left({{{\mathbb{E}}_{p\left(s\right)p\left(\hat{s}\right)}}[{e^{\mathcal{T}}}]}\right).$
@@
Note that $\mathcal{T}$ can be any class of functions satisfying integrability, as a result, the NN in $\mathcal{F}$ is employed as $\mathcal{T}$ to obtain a precise bound according to Eq. (6). Finally, the estimated mutual information can be denoted as
@@
[…]
Figure 3: The structure of the FDE network.
By optimizing the parameters through maximizing the mutual information, the first three dense layers in $\mathcal{F}$ learn to make the received feature vector approach the distributions of the original undistorted vectors across all the categories in the dataset. On this basis, the output vector $\bar{\bm{d}}^{j}$ as shown in Fig. 3 can serve as a weight vector that prominently quantifies the contributions of different received features to the information recovery, further reflecting the distortion levels of the input features after transmission [41, 42]. Specifically, as illustrated in [40], based on the chain rule and the smoothness of the estimated mutual information function, the correlation between the $i$-th output feature $\bar{{d}}^{j}_{i}$ of neural network $\mathcal{F}$ and the loss $L_{\text{FDE}}$ can be derived as follows,
$\Delta L_{\text{FDE}}(\bar{{d}}^{j}_{i})\approx\nabla_{\bar{{d}}^{j}_{i}}L_{\text{FDE}}\cdot\Delta\bar{{d}}^{j}_{i},$ (8)
where $\nabla_{\bar{{d}}^{j}_{i}}L_{\text{FDE}}$ represents the gradient of the loss function with respect to the output feature, $\Delta L_{\text{FDE}}(\bar{{d}}^{j}_{i})$ and $\Delta\bar{{d}}^{j}_{i}$ are the variation of the loss function and the output feature after iterations, respectively. Additionally, the output feature variation $\Delta\bar{{d}}^{j}_{i}$ is linearly related to its corresponding gradient, which can be represented as
$\Delta\bar{{d}}^{j}_{i}=-\eta\nabla_{\bar{{d}}^{j}_{i}}L_{\text{FDE}},$ (9)
where $\eta$ is the learning rate. Therefore, the variation of $L_{\text{FDE}}$ can be also represented as follows,
$\Delta L_{\text{FDE}}(\bar{{d}}^{j}_{i})\approx-\eta\,\bigl{\|}\nabla_{\bar{{d}}^{j}_{i}}L_{\text{FDE}}\bigr{\|}^{2}.$ (10)
This suggests that if $\bigl{|}\nabla_{\bar{{d}}^{j}_{i}}L_{\text{FDE}}\bigr{|}$ is larger, the loss decreases more rapidly, indicating that this feature contributes more to the reduction of the loss. As illustrated in Eq. (9), in the gradient descent process, the optimizer is more inclined to adjust these features, causing their variations to be more rapid to accelerate the reduction of the loss [43]. Finally, the output feature vector of $\mathcal{F}$ can display a more pronounced distribution corresponding to the decrease in loss (i.e., the increase in mutual information), which can be used to guide the feature distortion evaluation.
During testing, since the contributions of different received features to the information recovery are well-learned on all the categories, for an received feature vector $\hat{\bm{z}}_{\mathrm{p}}^{j}$ to feed into FDE, the output $\bar{\bm{d}}^{j}$ can represent the semantic similarity of $\hat{\bm{z}}_{\mathrm{p}}^{j}$ relative to the undistorted vectors that the model have learned corresponding to the specific category. Consequently, the distortion levels of features in received feature vector $\hat{\bm{z}}_{\mathrm{p}}^{j}$ can be inferred from output vector $\bar{\bm{d}}^{j}$ of FDE directly. Particularly, after applying the activation function to $\bar{\bm{d}}^{j}$, larger values indicate greater similarities, implying less distortion of the corresponding features in $\hat{\bm{z}}_{\mathrm{p}}^{j}$. Finally, these values form the vector ${\widetilde{\bm{{d}}}}^{j}$ by selecting the corresponding positions according to $\hat{\bm{z}}^{j}$, and the distortion measure vector ${\bm{{d}}}^{j}$ can be given by ${\bm{{d}}}^{j}=1-{\widetilde{\bm{{d}}}}^{j}$.
@@
Furthermore, to instruct the transmitter to retransmit the corrupted features, ${\bm{{d}}}^{j}$ is quantized into a binary feedback vector ${\bm{p}}^{(j+1)}$ to trigger the $(j+1)$-th retransmission. Note that the process is not trivial because the quantization is discrete, which makes the network non-differentiable and hard to optimize with back-propagation. Thus, a straight-through Gumbel-softmax estimator is investigated [44], where ${\bm{{d}}}^{j}$ is discretized in the forward pass but a gradient estimator based on Gumbel-softmax distributions is employed to resolve the non-differentiability in the backward pass. Specifically, for the forward propagation, ${\bm{p}}^{(j+1)}$ is drawn from ${\bm{{d}}}^{j}$ using a distortion threshold $t$, which can be represented as
${p}^{(j+1)}_{i}=\left\{\begin{array}[]{ll}1&\text{ if }{d}^{j}_{i}\geq t,\\ 0&\text{ otherwise, }\end{array}\right$. (11)
@@
where ${p}^{(j+1)}_{i}$ and ${d}^{j}_{i}$ denote the $i$-th value of ${\bm{p}}^{(j+1)}$ and $\bm{{d}}^{j}$}, respectively. On this basis, ${\bm{p}}^{(j+1)}$ can serve as a feedback vector to guide the next retransmission. Particularly, the number of features requested to be retransmitted is equal to the number of $1$ values in ${\bm{p}}^{(j+1)}$, which is given by
$R=\sum_{i=1}^{B}{p}^{(j+1)}_{i}$, (12)
where $B$ is the size of ${\bm{p}}^{(j+1)}$.